determine the amount of an investment if 700$ is invested at an interest rate of 8% compounded monthly for 9 years
Answer by stanbon(75887)

determine the amount of an investment if 700$ is invested at an interest rate of 8% compounded monthly for 9 years
-----------
Formula: A(t)=P(1+r/n)^(nt)
A(9) = 700(1+0.08/12)^(12*9)
A(9) = 700(1+0.00667)^(108)
A(9) = 700(2.05026
A(9) = $1435.18
===========
Cheers,
Stan H.